You are a landscape artist designing a patio. The square patio floor is to be made from 128 cubic feet of concrete. The thickness of the floor is 15.5 feet less than each side length of the patio. What are the dimensions of the patio floor?
Answer by lwsshak3(11628)

You are a landscape artist designing a patio. The square patio floor is to be made from 128 cubic feet of concrete. The thickness of the floor is 15.5 feet less than each side length of the patio. What are the dimensions of the patio floor
**
let x=each side length of square patio floor in feet
x-15.5=thickness of floor in feet
Volume of concrete floor=area*thickness
x^2(x-15.5)=128
x^3-15x^2-128=0
I don't know to solve this equation algebraically, but my computer graphing program gave an answer of x≈15.5307
ans:
side length of square patio floor≈15.5307 ft
thickness of the floor≈x-15≈.5307 ft